OpenStudy (thatonegirl_):

Probability !!! What is the probability of k>or=10, for 15 trials and .75 success probability?

OpenStudy (dumbcow):

For each k value: \[P(k) = \left(\begin{matrix}15 \\ k\end{matrix}\right)(.75)^k (.25)^{15-k}\] do it for k = 10,11,12,13,14,15 add all the probabilities to get final answer

OpenStudy (thatonegirl_):

There's no shorter way? My teacher mentioned something about the binomcdf function on the ti-84 calc.. I tried it but how do you specify whether it is greater than, less than, or equal to? @dumbcow

OpenStudy (thatonegirl_):

@Directrix